OpenStudy (anonymous):

A pool Is filled at a rate of 105 gal/ min what id the rate in ml per second

OpenStudy (wolf1728):

https://en.wikipedia.org/wiki/Gallon US customary liquid gallon (≈ 3.79 L) = 1.75 gallons/sec 1 gallon= 3,790 mL So, 105 gallons/minute = 6,632.50 mL / second

OpenStudy (jack1):

\[\frac{ 105 gal }{ 1\min }\times\frac{ 1\min }{60\sec }\times\frac{3 785.42mL }{ 1gal }=\frac{ x(mL) }{ 1\sec }\]